# Kuro Siwo: 33 billion under the water A global multi-temporal satellite dataset for rapid flood mapping

### v2

@@ 5 Discussion
Kuro Siwo evaluation: Quantitative comparison of Kuro Siwo with other datasets is challenging, as most publicly available SAR-based datasets do not follow the same evaluation principles. These principles include testing on events absent from the training set and ensuring a diverse array of test events.
We identify MMFlood as the most fitting dataset for our comparison, offering 34 GRD test events with labels acquired from CEMS. The authors report a maximum of mIoU solving the binary flood/no flood task, which is significantly lower than the one presented in BlackBench, where we solve a multi-class problem. This may be an indirect consequence of the problematic nature of CEMS labels as discussed in Sec. 3 and an indicator for the importance of accurate and cross verified ground truth labels. Another likely reason could be the oversight of the temporal dimension, as only post-event images are being used. Examining Tab.Ā 2, we observe that the temporal aspect is crucial for semantic segmentation models, with most performing best when utilizing all available pre-event images. Conversely, incorporating DEM information offers negligible improvements. Even when including DEM yields the best results for a model, the performance gains are marginal. Interestingly, we notice a consistent discrepancy across models in Tab.Ā 2 between the reported metrics for permanent water bodies and the rest of the classes. We hypothesize that this behaviour stems from the inherent challenge of discerning permanent water bodies from flooded areas in specific locations, especially close to overflowed rivers and lakes.
@@
This problem is magnified by the constant flow of water, resulting in small changes in the area it populates which in turn materializes with different signatures for each SAR timestep. This effect has a stronger presence in Fig. 3, where we present the per-event performance of the best model in BlackBench i.e. UNet-ResNet50. Overall, we notice a stable performance despite the variable environmental conditions between events. However, performance in permanent waters detection drastically decreases for Honduras and Australia. We provide a qualitative examination of samples from these areas in Fig. 4. Indeed, we observe good detection of both flooded areas and permanent water bodies, however, it is particularly challenging to accurately delineate their boundaries when they overlap. Notably, FloodViT performs comparably with most models and surpasses all change detection methods despite having only a small decoder trained for the segmentation task. We did not observe additional benefits from fine-tuning FloodViT-FT. Further investigation on SSL methods for this task is left as future work.
@@
SLC dataset: The inclusion of the unrefined SLC products paired with quality annotations at a global scale is an important aspect of Kuro Siwo. This component frees the end user of the task-specific preprocessing choices of the dataset creators, while enabling the investigation of important research questions such as the exploitation of the complex valued signal nature of SAR data. Our initial experiments suggest that DL models trained on SLC data are capable classifiers even on such unrefined data. In fact, a simple UNet with a ResNet18 backbone, trained on a subset of Kuro Siwo SLC dataset, is able to achieve F-Score on the binary water detection task and and on the flood and permanent waters categories respectively. We leave as future work the thorough exploration of the complex information of SAR data.
Limitations: Despite significant efforts to create a quality and diverse flood mapping dataset, approximately half of Kuro Siwoās flood events are from Europe, with the rest distributed globally, indicating a need for more balanced spatial coverage. Furthermore, while Kuro Siwo can train models to map observed flood extents from SAR images, accurately capturing the actual maximum flood extent remains dependent on the revisit time of the Sentinel-1 mission. Finally, our dataset and benchmark are based on a single modality C-band carrier frequency data, so it remains unclear how these models would generalize to X- and L-band data.

### v3

@@ 5 Discussion
Kuro Siwo evaluation: Quantitative comparison of Kuro Siwo with other datasets is challenging,
as most publicly available SAR-based datasets do not follow the same evaluation principles. These principles include testing on events absent from the training set and ensuring a diverse array of test events.
We identify MMFlood as the most fitting dataset for our comparison, offering 34 GRD test events with labels acquired from CEMS. The authors report a maximum of mIoU solving the binary flood/no flood task, which is significantly lower than the one presented in BlackBench, where we solve a multi-class problem. This may be an indirect consequence of the problematic nature of CEMS labels as discussed in Sec. 3 and an indicator for the importance of accurate and cross verified ground truth labels. Another likely reason could be the oversight of the temporal dimension, as only post-event images are being used. Examining Tab.Β 2, we observe that the temporal aspect is crucial for semantic segmentation models, with most performing best when utilizing all available pre-event images. Conversely, incorporating DEM information offers negligible improvements. Even when including DEM yields the best results for a model, the performance gains are marginal. Interestingly, we notice a consistent discrepancy across models in Tab.Β 2 between the reported metrics for permanent water bodies and the rest of the classes. We hypothesize that this behaviour stems from the inherent challenge of discerning permanent water bodies from flooded areas in specific locations, especially close to overflowed rivers and lakes.
@@
This problem is magnified by the constant flow of water, resulting in small changes in the area it populates which in turn materializes with different signatures for each SAR timestep. This effect has a stronger presence in Fig. 3, where we present the per-event performance of the best model in BlackBench, i.e. UNet-ResNet50. Overall, we notice a stable performance despite the variable environmental conditions between events. However, performance in permanent waters detection drastically decreases for Honduras and Australia. We provide a qualitative examination of samples from these areas in Fig. 4. Indeed, we observe good detection of both flooded areas and permanent water bodies, however it is particularly challenging to accurately delineate their boundaries when they overlap.
Notably, FloodViT performs comparably with most models and surpasses all change detection methods despite having only a small decoder trained for the segmentation task. We did not observe additional benefits from fine-tuning FloodViT-FT. Further investigation on SSL methods for this task is left as future work.
@@
SLC dataset: The inclusion of the unrefined SLC products paired with quality annotations at a global scale is an important aspect of Kuro Siwo. This component frees the end user of the task-specific preprocessing choices of the dataset creators, while enabling the investigation of important research questions such as the exploitation of the complex valued signal nature of SAR data. Our initial experiments (see Tab.2 in Supplementary material) suggest that DL models trained on SLC data are capable classifiers even on such unrefined data. In fact, a simple UNet with a ResNet18 backbone is able to achieve F-Score on the binary water detection task and and on the flood and permanent waters categories respectively. This is particularly promising since models in BlackBench are not tailored to the unique characteristics of SLC data. The thorough exploration of the complex information of SAR data is left as future work.
Limitations: Despite significant efforts to create a quality and diverse flood mapping dataset, approximately half of Kuro Siwoβs flood events are from Europe, with the rest distributed globally, indicating a need for more balanced spatial coverage. Furthermore, while Kuro Siwo can train models to map observed flood extents from SAR images, accurately capturing the actual maximum flood extent remains dependent on the revisit time of the Sentinel-1 mission. Finally, the inherent challenges associated with SAR data are a key limitation of Kuro Siwo. For example, speckle effect introduces granular noise that complicates the differentiation between flooded and non-flooded areas. Complex terrain and varying land cover types, such as forests, further complicate flood detection by affecting radar signal interaction. Variations in water surface roughness due to wind or the existence of vegetation can alter backscatter, making it difficult to consistently identify floodwaters. Detecting floods beneath dense vegetation is also challenging, as the S1 C-band radar signal may not penetrate sufficiently (as opposed to L-band SAR sensors) to reveal underlying water. Urban areas pose another difficulty, as radar signals often produce complex scattering effects (e.g. double-bounce reflections), making it difficult to distinguish between water and other surfaces, like wet roads or buildings.
Future research directions: Multispectral data, such as those provided by the Sentinel-2 constellation, can significantly aid flood mapping under cloud-free conditions since they provide a clearer view of the underlying waters compared to SAR. Incorporating multispectral imagery in Kuro Siwo would require extensive photointerpretation, as acquisition timestamps between the satellites would possibly differ leading to discrepancies in detected flood events, given the short life cycle of flash floods. Nevertheless, the synergistic use of both datasets is a promising research direction, combining the precision of multispectral data with the all-weather resilience of SAR. Having Kuro Siwo as a publicly available resource with reliable SAR based annotations can provide the foundation for assembling such a multi-modal dataset.
Moreover, including a permanent water layer could be beneficial for discriminating between the two water classes. However, water bodies exhibit dynamic variable extents due to a variety of factors including meteorological conditions and environmental changes. Using a static water layer as input to the models could potentially impute noise and ultimately disrupt the learning process. Refining existing layers for each timestep in Kuro Siwo could prove counterintuitive in production as rapid flood mapping is of utmost importance. Further investigation on such auxiliary information is left as future work.
Finally, a related but fundamentally different topic than rapid flood mapping is flood forecasting. Kuro Siwo is focused on supporting emergency response activities within short time frames rather than long-term monitoring of water bodies for disaster resilience. While crucial, this task demands a distinct approach and additional data, such as local weather patterns, water storage capacity, soil data, and denser time series. Future research on this topic could strongly complement Kuro Siwo.
@@ 6 Conclusion
Acknowledgments and Disclosure of Funding
This work has received funding from the project ThinkingEarth (grant agreement No 101130544) and from the project MeDiTwin (grant agreement No 101159723) of the European Unionβs Horizon Europe research and innovation programme.
Supplemental material
1 Dataset
The total size of the compressed dataset is TB, with the GRD component (including the DEMs and metadata) taking GB and the SLC GB.
All code and data will be maintained at the projectβs repo.
2 Evaluation on a new, unseen event
In Fig. 5 we assess the performance of our best model, i.e. Unet-ResNet50, on the recent floods of Emiglia-Romana, Italy, which took place on May 2023. Our predictions were based on a post-event SAR image acquired on 22/05/2023, and two pre-event SAR images from 10/05/2023 and 28/04/2023. On the right hand side of Fig. 5 we overlay our predictions with a Sentinel-2 RGB image captured on 23/05/2023 for better inspection of the annotated areas. Examining both derived maps we clearly see the capacity of the proposed model, showcasing the quality of Kuro Siwo. Our model manages to uncover all major water bodies, such as the rivers and ponds in the Pialassa della Baiona, as well as flooded regions of diverse sizes. Given that BlackBench presents baseline models, the performance of future task-specific methods may potentially surpass the baselines offered in this work.
3 Additional results
Fig. 6 showcases representative samples from the Kuro Siwo GRD dataset, providing a closer examination of flood events spanning a diverse set of countries. The predictions come from the best performing model in BlackBench, i.e UNet-ResNet50. The results underscore the significance of leveraging both polarization bands (VV and VH) for an accurate evaluation of the influence of the underlying land cover. While the VV band is effective in detecting the majority of flooded areas, the post-flood VV imagery may also include some bright regions in flooded areas that might be mistaken for dry land. These areas are refined and clarified by incorporating the VH band, which exhibits a more uniform dark region with smaller bright patches. For example, the flooded area in Fig. 6(e) is more concretely depicted in the VH band with stronger backscatter in the regions farthest from the river which also present a challenge to our model.
On the contrary, Fig. 6(b), Fig. 6(g) and Fig. 6(e) feature specific dark regions in the pre-flood imagery that are not labeled as water bodies in the annotation. These areas may correspond to isolated pixels experiencing the speckle effects or to crop parcels or vegetated areas with varying moisture content, that are in different phenological states compared to neighboring regions. In general, the model manages to successfully capture the outline of most water elements, albeit often missing narrow rivers, which is a rather difficult task considering the dynamic nature of water flow and the changes of river width over time.
Furthermore, Fig. 6(c) and Fig. 6(d) present instances of river overflow. The Kuro Siwo annotation precisely delineates the initial river shape as inferred by the two pre-event captions, along with the flooded river banks marked in purple. In Fig. 6(d), the Unet-ResNet50 achieves high accuracy in both permanent waters and flood, whereas in Fig. 6(c) it seems to miss the exact shape of the river in which is depicted in a more solid black region in the VH band. Furthermore, the overflowed river banks present a significant challenge for the model which is constantly confused between the two water classes.
Finally, Fig. 6(h) presents a particularly challenging sample, wherein the small reservoir of permanent water in the upper-left of the patch is difficult to detect due to the strong presence of SAR speckle noise.
4 Annotation process
The outlined annotation process unfolds in distinct phases: initially, we establish the entire procedure by defining the annotation categories, i.e. βPermanent Watersβ, βFloodsβ and βNo Waterβ. We extract the corresponding photointerpretation keys, set the annotation scale to 1:1000, and establish criteria for flagging artifacts in the input data. Next, we partition the flood events into five segments. A team of five SAR experts undertook the annotation process, with each member assigned to annotate one part and quality-check the annotations of another team member. The supervision of this process was overseen by a senior remote sensing scientist. Regular group meetings were conducted to address and resolve any potential discrepancies between the primary annotator and the responsible for the quality check.
5 Training setup
In Table 4 we report the training hyperparameters for each model in BlackBench. All SAR images were clipped at a max value of 0.15 and normalized to 0 mean and 1 standard deviation. At training time we skip samples that do not contain water (regardless of the class). We test on all available samples of the test areas. All models utilize the cross-entropy loss function, apart from SNUNet-CD that employs a compound loss consisting of cross-entropy and Dice loss, in line with the original publication.
6 SLC benchmark
In Table 3 we present the BlackBench results for the SLC component of Kuro Siwo.
UNet-ResNet18 | 3 | - | β | 97.88 | 76.76 | 71.20 | 71.14 | 79.94 |
UNet-ResNet50 | 3 | - | β | 97.68 | 78.76 | 68.91 | 71.00 | 78.10 |
UNet-ResNet101 | 3 | - | β | 97.68 | 76.64 | 69.74 | 70.38 | 77.81 |
DeepLab-ResNet18 | 3 | - | β | 97.63 | 74.21 | 69.15 | 69.08 | 78.51 |
DeepLab-ResNet50 | 2 | - | β | 97.63 | 76.57 | 66.20 | 68.96 | 77.10 |
DeepLab-ResNet101 | 3 | - | - | 97.63 | 74.92 | 69.72 | 69.60 | 77.13 |
UPerNet-SwinS | 3 | - | - | 97.71 | 79.48 | 70.39 | 71.93 | 78.17 |
UPerNet-SwinB | 3 | - | - | 97.63 | 79.16 | 70.25 | 71.68 | 77.23 |
UPerNet-ConvnextS | 3 | - | - | 97.64 | 78.93 | 69.33 | 71.22 | 77.45 |
UPerNet-ConvnextB | 3 | - | - | 97.71 | 78.79 | 69.70 | 71.34 | 77.51 |
FloodViT | 3 | - | - | 97.32 | 60.03 | 56.87 | 59.13 | 69.76 |
FloodViT-FT | 3 | - | - | 96.84 | 51.23 | 57.62 | 56.26 | 61.43 |
FC-EF-Diff | 2 | β | - | 97.53 | 70.51 | 69.13 | 67.49 | 75.97 |
FC-EF-Conc | 2 | - | - | 97.72 | 71.92 | 68.78 | 68.03 | 77.3 |
SNUNet-CD | 2 | - | β | 97.65 | 75.31 | 68.82 | 69.42 | 77.99 |
Changeformer | 2 | β | - | 97.61 | 69.31 | 69.34 | 67.14 | 75.87 |
ConvLSTM | 3 | - | - | 97.76 | 78.16 | 70.58 | 71.44 | 78.0 |
7 Computational resources
All experiments were conducted on a single NVIDIA GeForce RTX 3090 Ti.
8 Intended uses
Kuro Siwo is designed to address two primary objectives: a) overcoming the lack of large, meticulously annotated datasets for flash flood mapping and b) promoting the development of deep learning methods that utilize both phase and amplitude information, without restrictive preprocessing choices. The intended uses of Kuro Siwo are thus, a) the development of robust and accurate flash flood mapping methods, which could potentially be applied in operational settings, and b) research in the direct exploitation of complex SAR data.
Model | Optimizer | LR | LR scheduling |
Weight
decay |
# epochs |
UNet-ResNet18 | Adam | 1e-3 | Cosine | - | 20 |
UNet-ResNet50 | Adam | 1e-3 | Cosine | - | 20 |
UNet-ResNet101 | Adam | 1e-3 | Cosine | - | 20 |
DeepLab-ResNet18 | Adam | 1e-3 | Cosine | - | 20 |
DeepLab-ResNet50 | Adam | 1e-3 | Cosine | - | 20 |
DeepLab-ResNet101 | Adam | 1e-3 | Cosine | - | 20 |
UPerNet-SwinS | Adam | 1e-4 | - | 1e-4 | 20 |
UPerNet-SwinB | Adam | 1e-4 | - | 1e-4 | 20 |
UPerNet-ConvnextS | Adam | 1e-4 | - | 1e-4 | 20 |
UPerNet-ConvnextB | Adam | 1e-4 | - | 1e-4 | 20 |
FloodViT | Adam | 1e-3 | Cosine | 1e-5 | 20 |
FloodViT-FT | Adam | 1e-4 | Cosine | 1e-3 | 20 |
FC-EF-Diff | Adam | 1e-5 | - | - | 150 |
FC-EF-Conc | Adam | 1e-5 | - | - | 150 |
SNUNet-CD | Adam | 1e-5 | - | - | 150 |
Changeformer | SGD | 6e-4 | Linear | 1e-5 | 150 |
ConvLSTM | Adam | 1e-3 | - | - | 150 |
9 License
The authors state that they bear all responsibility in case of any rights violation. The Kuro Siwo dataset as well as BlackBench are released under the MIT License.333https://opensource.org/license/MIT
10 Maintenance plan
The main hosting platform of Kuro Siwo and BlackBench is GitHub and the authors guarantee access to the main code, as well as maintenance and issue tracking. Due to their seer size, data are hosted at a major cloud storage provider, ensuring seamless availability and adequate download speeds.